OpenStudy (anonymous):

Solve x2 + 6x - 5 = 0 by completing the square. Which equation is used in the process?

OpenStudy (anonymous):

A trick on getting your last number (ie. x^2+6x+? )for a complete square is by dividing the middle number in half and then squaring it. So: 6/2=3....3^2=9 x^2+6x-5+14=14 (remember to always add/subtract on both sides of the equation) So: x^2+6x+9=14 And so: (x+3)*(x+3) = 14 or (x+3)^2 = 14